# Hongyi Du1, Jiarui Yao1, Xiaocheng Yang1, Denghui Zhang1,3, Yunzhu Li4, Heng Ji1 1

### v1

@@ Abstract
Language model agents excel in long-session planning and reasoning, but existing benchmarks primarily focus on goal-oriented tasks with explicit objectives, neglecting creative adaptation in unfamiliar environments. To address this, we introduce EscapeBench—a benchmark suite of room escape game environments designed to challenge agents with creative reasoning, unconventional tool use, and iterative problem-solving to uncover implicit goals. Our results show that current LM models, despite employing working memory and Chain-of-Thought reasoning, achieve only 15% average progress without hints, highlighting their limitations in creativity. To bridge this gap, we propose EscapeAgent, a framework designed to enhance creative reasoning through Foresight (innovative tool use) and Reflection (identifying unsolved tasks). Experiments show that EscapeAgent can execute action chains over 1,000 steps while maintaining logical coherence. It navigates and completes games with up to 40% fewer steps and hints, performs robustly across varying difficulty levels, and achieves higher action success rates with more efficient and innovative puzzle-solving strategies. All the data and codes are released [^1] https://github.com/qiancheng0/EscapeBench.
[…]
[Uncaptioned image] EscapeBench: Pushing Language Models to Think Outside the Box
@@ 1 Introduction
Building robust language model (LM) agents for planning and reasoning has always been challenging. Recent efforts have explored how agents could compress and utilize memory (Wang et al., 2023a; Hu et al., 2023; Liu et al., 2023b; Liang et al., 2023b; Wang et al., 2024d, c; Zhong et al., 2024), perform complex reasoning (Wei et al., 2022; Kojima et al., 2022; Zhou et al., 2023a; Lin et al., 2024; Yao et al., 2023), planning (Wang et al., 2023b; Liu et al., 2023a; Hao et al., 2023; Yao et al., 2024; Zhou et al., 2024a), and reflection (Madaan et al., 2024; Zhang et al., 2024a, b; Miao et al., 2024; Dhuliawala et al., 2024) to improve task success rate. Integrating these capabilities, recent lines of work begin to build agents for embodied actions (Zheng et al., 2024; Huang et al., 2024a; Zhu et al., 2023) and tool use  (Schick et al., 2023; Qin et al., 2023; Qian et al., 2024) grounded in environments including the Web (Nakano et al., 2021; Furuta et al., 2024; Gur et al., 2024), games (Guo et al., 2023; Xu et al., 2023; Hu et al., 2024), and society (Park et al., 2023; Liu et al., 2023c; Li et al., 2023; Ren et al., 2024).
@@
The surge of agent systems also accelerates the development of simulation environments to evaluate an agent’s capabilities, including tasks like computer-based operations (Yao et al., 2022; Deng et al., 2024; Zhou et al., 2024b; Xie et al., 2024; Liu et al., 2024b), scientific research (Wang et al., 2022; Bran et al., 2023; Boiko et al., 2023; Huang et al., 2023a), and interactive experiences in text-based (Côté et al., 2019; Urbanek et al., 2019; O’Gara, 2023; Wu et al., 2024) or virtual sandbox game environments (Lin et al., 2023; BAAI, 2023; Wang et al., 2024a). However, most existing benchmarks are goal-oriented, emphasizing models’ planning, reasoning, and error-handling abilities, while overlooking an agent’s creativity: the capacity to think innovatively and adapt its observations to new, emergent, and unstructured scenarios.
Current agents still significantly lack creativity in novel tool use (Zhang et al., 2023), as their training predominantly focuses on memorizing tool-task associations. This emphasis overshadows their ability to explore tool affordances and adapt to unstructured scenarios (Zhang et al., 2024c), both of which are critical for innovation. In robotics and autonomous systems, this limitation prevents agents from leveraging tools in unexpected ways to tackle unforeseen challenges (Bossema et al., 2023; Sandoval et al., 2022).
In fact, creativity is widely recognized as a crucial component of intelligence. In cognitive science, the well-established Triarchic Theory of Intelligence (Sternberg, 1984) divides intelligence into three components: practical, analytical, and creative. While current reasoning benchmarks primarily assess analytical intelligence through problem-solving, and simulation environments focus on practical intelligence by testing the application of knowledge in real-world scenarios, creative intelligence remains largely unaddressed. This gap is notable given its importance for enabling agents to demonstrate human-like adaptability and think beyond predefined parameters.
To bridge this gap, we introduce EscapeBench, a benchmark that evaluates LM’s creative reasoning using scenarios inspired by room escape games. These scenarios test whether a model can escape a room through trial and error by solving unconventional puzzles that require “thinking outside the box” skills including creative observation, innovative tool use, and strategic problem-solving. As shown in Figure 1, a wooden stick, typically used for walking or poking, has to be repurposed to pry open a lid due to its hard texture. This demands agents to perform adaptive reasoning under customized constraints. Overall, our benchmark sets itself apart through several distinctive features:
@@
We benchmarked multiple open- and closed-source models within the BaseAgent framework, which incorporates working memory and Chain-of-Thought reasoning (Wei et al., 2022), by objectively evaluating their performance based on the total steps and hints used to complete each game. Our results show that even the best models struggle to complete the easiest game setting without hints, often requiring up to ten times the optimal steps and falling far behind human performance. These findings highlight how models tend to be constrained by conventional thinking patterns, struggling to break free from them and show creativity.
To address this, we further introduce EscapeAgent, enhanced with Foresight for creative tool use and Reflection for implicit goal identification. The Foresight module encourages the agent to propose and self-evaluate creative hypotheses and predict outcomes for tool use before taking action, while the Reflection module helps the agent maintain an unsolved task list, providing dynamic reference to enhance its awareness of goals in future actions. Experiments show that EscapeAgent reduces hint reliance by nearly 50% and significantly lowers the number of action steps required to complete the game. It performs robustly across different difficulty levels, making progress more efficiently and exploring creative strategies with a higher success rate. In summary:
@@ 2 Related Work
Creativity is a cornerstone of human intelligence and a growing focus in AI research Legg and Hutter (2007); Lake et al. (2017). LMs have demonstrated notable creative capabilities across domains - they excel at generating narratives and poetry (Brown et al., 2020; Akoury et al., 2020), show effectiveness in tool creation and design (Qian et al., 2023; Cai et al., 2024), and augment human creativity through interactive ideation (Mialon et al., 2023). In scientific discovery, research has also found that LM-generated ideas tend to be more novel but slightly less feasible than those from human experts (Si et al., 2024; Wang et al., 2024b).
@@
Existing work in agent evaluation primarily focuses on text-based and sandbox-like environments that assess an agent’s cognitive and behavioral abilities in goal-oriented tasks Zhou et al. (2023b); Chen et al. (2022, 2024); Yu et al. (2024); Deng et al. (2024), with emerging research exploring the integration of LM/VLM-enabled agents in robotics to address challenges in real-world, embodied settings Liang et al. (2023a); Huang et al. (2023b, 2024b); Rana et al. (2023). Text-based environments Yuan et al. (2018); Côté et al. (2019), such as interactive fiction games Lin et al. (2024) or conversational agents Qiao et al. (2023), evaluate natural language understanding, reasoning, and decision-making consistency Uludağlı and Oğuz (2023); Qi et al. (2024). Games like Zork Infocom (1980) and TextWorld measure narrative comprehension and problem-solving in structured contexts. In contrast, sandbox environments Lin et al. (2023); Gan et al. (2021); Fan et al. (2022) like Minecraft Zhu et al. (2023) and Roblox Rospigliosi (2022) provide more open-ended settings that test spatial reasoning, planning, and collaboration Carroll et al. (2019); Agashe et al. (2023). Evaluation of agents in these environments typically relies on task-specific, cognitive, behavioral, and learning metrics to assess their ability to achieve objectives, reason adaptively, and collaborate. However, existing benchmarks often neglect creative adaptation in unfamiliar environments requiring unconventional tool use and proactive problem-solving. To address this gap, we introduce EscapeBench which evaluates agents’ creative reasoning and strategic problem-solving through navigating uncertain goal-achieving pathways, offering a new dimension of agent evaluation.
Figure 2: An illustration of Scenes, Tools, and Items in the game engine design, and their relations with the model agent’s action space. Tools can be collected and used to craft new tools or applied to other items, while an item typically requires a click, input, or tool application to trigger specific effects.
@@ 3 EscapeBench Construction
3.1 Motivation
Most current agent benchmarks emphasize explicit, goal-oriented tasks, where decisions rely on commonsense knowledge encoded within the agent model’s parametric space. For instance, a typical task in AgentBench (Liu et al., 2024b) might ask: “What musical instruments did the Minnesota-born Nobel prize winner play?” Using the Freebase search API, the agent can chart a clear pathway to the goal, identifying relevant data and filtering out irrelevant results. This process is deterministic and fact-based, with each step reflecting the agent’s analytical skill in processing search outputs. While such benchmarks effectively test an agent’s information-seeking and retrieval capabilities, they often overlook its ability to engage in creative problem-solving—an area that extends beyond analytical and practical intelligence.
This leads to our core research question: How to build an environment that benchmarks an agent’s creative intelligence? To answer this, an agent must demonstrate out-of-the-box thinking in its reasoning. Since the primary function of agents lies in their tool-use abilities, we aim to design a benchmarking environment that emphasizes creative tool use. Room escape game scenarios, by nature, are well-suited for this purpose, as they require a player to creatively make use of all the things at hand to solve difficult puzzles. In the following sections, we introduce the components of our game engine (Section 3.2), the action space available to agents (Section 3.3), the data annotation process (Section 3.4), and preliminary case study results (Section 3.5).
@@ 3.2 Engine Design
Our engine aims to simulate the room escape environment that i) receives agent actions and ii) makes corresponding environment responses as agent action’s reward. It builds up a logical skeleton that ensures a smooth benchmarking process once the game configuration is given. Specifically, our game engine involves three key components:
@@
The interaction of these three key components forms a game scenario’s basic logic. For the game engine example in Figure 2, different scenes inside a van connect with each other forming a graph to represent their physical connectivity via connectors, such as doors or tunnels. The tool key chain collected in a scene will be placed in the bag for future use, while the item wire iron is waiting for something sharp to cut it off to trigger some effects. Please refer to Appendix A for more detailed examples and explanations.
@@ 3.3 Action Space
The model agent could take five different actions in each scene. While the action space is well-defined, the parameter space—regarding the scenes, items, or tools involved in these actions—is high-dimensional, thus enabling dynamic interactions.
@@
Figure 2 illustrates the connections between the components of the game engine and the model agent’s action space. Among all the actions, “Apply” and “Craft” stand out as the most creativity-driven, as they require the agent to think innovatively about how to use or craft tools in an unseen way during its training. We delve into specific examples in Section 3.5.
@@ 3.4 Annotation and Statistics
Building on existing online room escape games and puzzle-solving logic [^2] https://spotlight.ee, we present EscapeBench, featuring 12 scenarios and 36 settings across three difficulty levels. Each scenario, such as a laundry room or basement exploration, is manually annotated for high-quality game engine performance. Difficulty levels are defined by the clarity of descriptions and the feedback provided for unexpected actions, detailed in Table 1. These scenarios emphasize creative tool use and crafting strategies, challenging agents throughout the game, making EscapeBench a robust environment for testing creativity. We make all our scenarios text-based to accommodate the evaluation of top-performing language models. We present in Figure 3 detailed statistics on total scenes, tools, items, and key steps involved in all game settings.
@@ 3.5 Preliminary Study
To verify the creative reasoning required in our designed games and assess the current LMs’ ability to perform such reasoning, we conduct a preliminary case study in Table 2. We sample diverse real scenarios from EscapeBench’s game design and test them on closed-source GPT-4o (Hurst et al., 2024) and open-source LLama-3.1-70B (Dubey et al., 2024). For each model, we instruct it to think creatively about what it might achieve with the provided Items and Tools in room escape scenarios.
Our results reveal that: i) EscapeBench offers a wide range of creative reasoning challenges, encompassing unconventional tool usage, implicit numerical puzzle solving, innovative tool crafting strategies, etc. ii) Both closed- and open-source models demonstrate significant limitations in creativity and imagination, particularly in identifying implicit goals and devising creative strategies. These findings emphasize the complexity and necessity of our benchmark as well as the current gaps in models’ creativity.
@@ 4 EscapeAgent Design
Our preliminary study highlights that current LMs often struggle with identifying implicit goals and engaging in the creative reasoning needed to solve complex puzzles. To overcome these challenges and maximize the LM’s potential, we present EscapeAgent, a novel framework designed to uncover tasks dynamically and reason creatively about tool use. EscapeAgent addresses two core challenges identified by EscapeBench:
  • Uncertain Goal Pathways: We introduce a Reflection module, which dynamically maintains a task list by adding, updating, or removing tasks after each trial-and-error action. This approach fosters proactive task discovery and sharpens the agent’s focus on specific goals (Section 4.2).
  • Creative Tool Use: We design a Foresight module that facilitates explicit reasoning about creative tool applications and their alignment with specific tasks. It enables the agent to hypothesize and evaluate strategies before execution, thus promoting problem-solving (Section 4.3).
By integrating both modules with a BaseAgent, EscapeAgent excels at managing Super-long Reasoning Chains, effectively performing over 1,000 consecutive actions while maintaining logical coherence. This robust design greatly enhances the language model’s capabilities in creativity, problem-solving, and strategic thinking.
@@ 4.2 Reflection Module
The Reflection Module enables the agent to maintain a structured task list, updated through:
@@
Each task list entry includes the task name, the target item, and previously failed actions. This organization prevents the agent from repeating ineffective actions, thus promoting efficiency.
The reflection module is triggered after every non-move action, using environment feedback to guide task list updates. For example, in Figure 4 (right), Task 1 is deleted once the machine successfully starts. By encouraging targeted problem-solving rather than random exploration, the reflection module significantly improves task-solving efficiency. See Section C.2 for more details.
@@ 4.3 Foresight Module
  • When a new task is identified: The agent evaluates existing tools it has and hypothesizes potential actions to achieve the task.
  • When a new tool is collected: The agent assesses how the tool might contribute to solving existing tasks or crafting new tools.
If the module generates at least one valid action hypothesis, the agent enters a “Try Action” state, where it systematically tests the proposed actions in subsequent steps. Otherwise, it remains in a “Free Explore” state, operating like the BaseAgent to explore the environment.
For instance, in Figure 4, the foresight module identifies the action of clicking a button on the machine, thus guiding the agent into the “Try Action” state. Overall, this module enables the agent to adapt flexibly under customized scenarios, make bold hypotheses, and execute targeted action trials efficiently. See Section C.3 for further details.
5 Experiments
We divide our experiments into two parts: i) Benchmarking the creativity of current models using the BaseAgent (Section 5.2), and ii) Demonstrating the effectiveness of EscapeAgent (Section 5.3).
5.1 Settings
Environment. Experiments are conducted on 36 game settings. An agent is considered to be making progress if it either achieves a key step (bottleneck actions defined in Figure 3) or collects a new tool. Agents will receive help if they fail to make progress for 50 consecutive actions (see Section D.1), thus ensuring full completion of the game. The working memory length is set to 10 due to context window limitations.
Models. We evaluate both closed- and open-source models: GPT-4o, GPT-4o-mini (Hurst et al., 2024), Claude-3.5 (Anthropic, 2024), Gemini-1.5 (Team et al., 2024), Llama-3.1 (Dubey et al., 2024), Qwen-2.5 (Team, 2024), DeepSeek-LLM (Liu et al., 2024a), Phi-3.5 (Abdin et al., 2024), Yi (Young et al., 2024), Ministral (MistralAI, 2024). Models with fewer than 7B parameters are excluded due to near-random behavior. For consistency, we set the sampling temperature to $T=0$ and $n=1$.
Metrics. We use two main metrics including:
  • Hints Used: Total hints used in a game.
  • Total Steps: Total actions taken in a game.
Auxiliary metrics for analysis include:
Results are micro-averaged across the 36 settings.

### v2

# Language Model Agents

@@ Abstract
Language model agents excel in long-session planning and reasoning, but existing benchmarks primarily focus on goal-oriented tasks with explicit objectives, neglecting creative adaptation in unfamiliar environments. To address this, we introduce EscapeBench—a benchmark suite of room escape game environments designed to challenge agents with creative reasoning, unconventional tool use, and iterative problem-solving to uncover implicit goals. Our results show that current LM models, despite employing working memory and Chain-of-Thought reasoning, achieve only 15% average progress without hints, highlighting their limitations in creativity. To bridge this gap, we propose EscapeAgent, a framework designed to enhance creative reasoning through Foresight (innovative tool use) and Reflection (identifying unsolved tasks). Experiments show that EscapeAgent can execute action chains over 1,000 steps while maintaining logical coherence. It navigates and completes games with up to 40% fewer steps and hints, performs robustly across difficulty levels, and achieves higher action success rates with more efficient and innovative puzzle-solving strategies. All the data and codes are released [^1] https://github.com/qiancheng0/EscapeBench.
[Uncaptioned image] EscapeBench: Towards Advancing Creative Intelligence of
Language Model Agents
@@ 1 Introduction
Building robust language model (LM) agents to perform planning and reasoning has always been a challenging task. Recent efforts have explored how agents could compress and utilize memory (Wang et al., 2023a; Hu et al., 2023; Liu et al., 2023b; Liang et al., 2023b; Wang et al., 2024c; Zhong et al., 2024), perform complex reasoning (Wei et al., 2022; Kojima et al., 2022; Zhou et al., 2023a; Lin et al., 2024; Yao et al., 2023), planning (Wang et al., 2023b; Liu et al., 2023a; Hao et al., 2023; Yao et al., 2024; Zhou et al., 2024a), and reflection (Madaan et al., 2024; Zhang et al., 2024a, b; Miao et al., 2024; Dhuliawala et al., 2024) to improve task success rate. Integrating these capabilities, recent lines of work begin to build agents for embodied actions (Zheng et al., 2024; Huang et al., 2024a; Zhu et al., 2023) and tool use  (Schick et al., 2023; Qin et al., 2023; Qian et al., 2024; Wang et al., 2025; Qian et al., 2025a) grounded in environments including the Web (Nakano et al., 2021; Furuta et al., 2024; Gur et al., 2024), games (Guo et al., 2023; Xu et al., 2023; Hu et al., 2024), and society (Park et al., 2023; Liu et al., 2023c; Li et al., 2023; Ren et al., 2024).
@@
The surge of LM agent systems also accelerates the development of simulation environments, including tasks like computer-based operations (Yao et al., 2022; Deng et al., 2024; Zhou et al., 2024b; Xie et al., 2024; Liu et al., 2024b), scientific research (Wang et al., 2022; Bran et al., 2023; Boiko et al., 2023; Huang et al., 2023a), and interactive experiences in text-based (Côté et al., 2019; Urbanek et al., 2019; O’Gara, 2023; Wu et al., 2024) or virtual sandbox game environments (Lin et al., 2023; BAAI, 2023; Wang et al., 2024a). However, most existing benchmarks are usually goal-oriented, emphasizing models’ planning, reasoning, and error-handling abilities, while overlooking their creativity: the capacity to think innovatively and adapt their observations to new, unstructured scenarios.
Current agents still significantly lack creativity in novel tool use (Zhang et al., 2023), as their training predominantly focuses on memorizing tool-task associations. This emphasis overshadows their ability to explore tool affordances and adapt to unstructured scenarios (Zhang et al., 2024c). Despite this, creativity is still widely recognized as a crucial component of intelligence. In cognitive science, the well-established Triarchic Theory of Intelligence (Sternberg, 1984) divides intelligence into three components: practical, analytical, and creative. While current reasoning benchmarks primarily assess analytical intelligence through problem-solving, and simulation environments focus on practical intelligence by testing knowledge application in real-world scenarios, creative intelligence remains largely unaddressed.
To bridge this gap, we introduce EscapeBench, a benchmark that evaluates LM’s creative reasoning using scenarios inspired by room escape games. These scenarios challenge conventional thinking through unusual settings and require “thinking outside the box” skills including creative tool use and strategic problem-solving. As shown in Figure 1, a wooden stick, typically used for walking or poking, has to be repurposed to pry open a lid due to its hard texture. This demands agents to perform adaptive reasoning under customized constraints. Overall, our benchmark has several distinctive features:
@@
We benchmarked multiple models within the BaseAgent framework, which incorporates working memory and Chain-of-Thought reasoning (Wei et al., 2022). Our results show that even the best models struggle to complete the easiest game setting without hints, often requiring up to ten times the optimal steps and falling far behind human performance. These findings highlight how models tend to be constrained by conventional thinking patterns, struggling to break free and show creativity.
To overcome this limitation, we introduce EscapeAgent, enhanced with Foresight for creative tool use and Reflection for implicit goal identification. Foresight enables the agent to propose and evaluate tool-use hypotheses before acting, while Reflection maintains an unsolved task list to guide future actions. Experiments show EscapeAgent reduces hint reliance by nearly 50%, lowers total action steps, and performs robustly across difficulty levels, achieving more efficient progress and higher action success rates with creative strategies. Our contributions include the following:
@@ 2 Related Work
Creativity is a cornerstone of human intelligence and a growing focus in AI research Legg and Hutter (2007); Lake et al. (2017). LMs have demonstrated notable creative capabilities across domains - they excel at generating narratives and poetry (Brown et al., 2020; Akoury et al., 2020), show effectiveness in tool creation and design (Qian et al., 2023; Cai et al., 2024), model real-world challenges (Qian et al., 2025b), and augment human creativity through interactive ideation (Mialon et al., 2023). In scientific discovery, research has also found that LM-generated ideas tend to be more novel but slightly less feasible than those from human experts (Si et al., 2024; Wang et al., 2024b).
@@
Agent evaluation often focuses on text-based or sandbox environments for assessing cognitive and behavioral abilities in goal-oriented tasks Zhou et al. (2023b); Chen et al. (2022, 2024); Yu et al. (2024); Deng et al. (2024), with emerging work exploring LM/VLM-enabled agents in robotics for real-world challenges Liang et al. (2023a); Huang et al. (2023b, 2024b); Rana et al. (2023); Zhu et al. (2025); Yang et al. (2025). Text-based environments Yuan et al. (2018); Côté et al. (2019), such as interactive fiction games Lin et al. (2024) or conversational agents Qiao et al. (2023), evaluate natural language understanding, reasoning, and decision-making consistency Uludağlı and Oğuz (2023); Qi et al. (2024). Games like Zork Infocom (1980) and TextWorld measure narrative comprehension and problem-solving in structured contexts. In contrast, sandbox environments Lin et al. (2023); Gan et al. (2021); Fan et al. (2022) like Minecraft Zhu et al. (2023) and Roblox Rospigliosi (2022) provide more open-ended settings that test spatial reasoning, planning, and collaboration Carroll et al. (2019); Agashe et al. (2023). The settings typically rely on task-specific metrics for goal achievement but overlook creative and proactive problem-solving in unfamiliar contexts. To address this, we introduce EscapeBench to evaluate agents’ creative reasoning in navigating uncertain goal pathways, offering a novel approach to agent assessment.
Figure 2: An illustration of Scenes, Tools, and Items in the game and their relations with agent action space. Tools can be collected for “Apply” and “Craft”, while items require “Input”, “Click” or “Apply” of tools to trigger effects.
@@ 3 EscapeBench Construction
Most agent benchmarks focus on explicit, goal-oriented tasks grounded in commonsense knowledge, where agents can chart clear pathways to achieve goals using analytical and practical intelligence, but they often overlook creative intelligence. This raises our core research question: How to build an environment that benchmarks an agent’s creative intelligence? Given that tool use is central to agent functionality, we propose room escape game scenarios, which naturally require creative tool use to solve complex puzzles, as an ideal environment for this evaluation.
@@ 3.1 Engine Design
Our game engine aims to simulate the room escape environment that i) receives agent actions and ii) makes corresponding environment feedback as agent action’s reward. Specifically, our game engine involves three key components:
@@
The interaction of these components defines a game’s basic logic. In the van example from Figure 2, scenes are connected into a graph, representing physical connectivity via doors or tunnels. The tool key chain is collected in the bag for future use, while the wire iron awaits something sharp to cut it to trigger effects. Please refer to Appendix A for more detailed examples and explanations.
@@ 3.2 Action Space
The model agent could take five different actions. While the action space is well-defined, the parameter space—regarding the scenes, items, or tools involved in these actions—is high-dimensional, thus allowing for dynamic interactions.
@@
Figure 2 illustrates the connections between game engine components and agent action space. Among all the actions, “Apply” and “Craft” stand out as the most creativity-driven, as they require the agent to think innovatively about how to use or craft tools in an unseen way during its training. We delve into specific examples in Section 3.4.
@@ 3.3 Annotation and Statistics
Building on existing online room escape games and puzzle-solving logic [^2] https://spotlight.ee, we introduce EscapeBench, featuring 36 game settings spanning three difficulty levels. Each scenario includes three versions with consistent solution logic but varying in description clarity and feedback granularity. A detailed example is shown in Table 1. All scenes, items, and tools are manually annotated to ensure high quality. These scenarios emphasize creative tool use and crafting strategies, challenging agents throughout the game, making EscapeBench a robust environment for testing creativity. A detailed statistic is presented in Figure 3.
@@ 3.4 Preliminary Study
We sample scenarios from EscapeBench and test GPT-4o (Hurst et al., 2024) and LLama-3.1-70B (Dubey et al., 2024)’s creative reasoning performance through case studies in Table 2. Our results reveal that: i) EscapeBench presents diverse creative reasoning challenges, including unconventional tool use, implicit numerical puzzles, and innovative tool crafting. ii) Both closed- and open-source models struggle with creativity, especially in identifying implicit goals and forming creative strategies. These findings highlight the complexity of EscapeBench and the gaps in model creativity.
@@ 4 EscapeAgent Design
To address challenges identified in the preliminary study, we introduce EscapeAgent, a framework addressing two core issues from EscapeBench:
  • Uncertain Goal Pathways: A Reflection module dynamically manages a task list, refining goals through trial-and-error to enhance action focus and proactive task discovery (Section 4.2).
  • Creative Tool Use: A Foresight module enabling explicit reasoning about tool applications, allowing the agent to hypothesize and evaluate strategies before execution (Section 4.3).
Integrating both modules with a BaseAgent, EscapeAgent excels in handling Super-long Reasoning Chains and significantly boosts the model’s creativity, problem-solving, and strategic thinking.
@@ 4.2 Reflection Module
The Reflection Module manages a structured task list updated through three actions:
@@
Each entry includes the task name, target item, and failed actions, preventing repeated mistakes and improving efficiency. Triggered after non-move actions, it uses feedback to update the task list. For example, in Figure 4 (right), Task 1 is deleted once the machine starts, encouraging focused problem-solving over random exploration. See Section C.2 for details.
@@ 4.3 Foresight Module
  • New Task Identified: The agent hypothesizes potential actions to achieve it using available tools.
  • New Tool Collected: The agent assesses its use for solving existing tasks or crafting new tools.
If a valid hypothesis is proposed, the agent enters “Try Action” state to test it; otherwise, it stays in “Free Explore” state, operating like the BaseAgent. For example in Figure 4 (left), the agent identifies clicking the button as an action worth trying, thus guiding it into the “Try Action” state. It enables the agent to adapt flexibly under customized scenarios, make bold hypotheses, and execute targeted trials efficiently. See Section C.2 for details.